# Perspective: Time irreversibility in systems observed at coarse resolution

### v2

@@ IV Outlook
Recently there has been a surge of interest in understanding and inferring irreversibility of coarse-grained observables. However, despite all efforts and advances the picture is far from complete and many fundamental (and occasionally elementary) questions remain elusive.
Semi-Markov processes and milestoning. First, while the importance of semi-Markov processes (of order $n$) in stochastic thermodynamics is visibly growing, it seems to be underappreciated that parameterizing these requires much more effort and statistics. For example, whereas a Markov process is fully specified with all transition rates between states, a first-order semi-Markov process (i.e. renewal dynamics) requires for each given state the respective splitting probabilities and waiting-time distributions to transition into each of the neighboring states, respectively (and the waiting-time distributions require a parameterization themselves, say a sum of exponential functions). What is worse, a second-order semi-Markov process requires for each given state the splitting probabilities and waiting-time distributions to transition into each of the neighboring states conditioned on all incoming transitions, (and the waiting-time distributions require a parameterization themselves). Typically we do not know a priori the order of the semi-Markov process unless we know both, the microscopic dynamics as well as the coarse graining. Moreover, the parameterization of semi-Markov processes becomes exponentially more challenging with the semi-Markov order. From a practical point of view one should therefore strive to approach renewal dynamics and hence invoke milestoning. However, formally milestoning has not yet been introduced to stochastic thermodynamics beyond renewal processes [52]. Moreover, we lack (sharp) thermodynamic inequalities, such as thermodynamic uncertainty relations and transport bounds, for milestoned dynamics. Another intriguing question, inspired by Ref. 57, concerns thermodynamic inference via milestoning of milestoned trajectories.
Thermodynamics of semi-Markov processes. Continuing with semi-Markov processes, rigorous results are only available for renewal dynamics. The results for the thermodynamic entropy production rate for general $n\geq 2$ semi-Markov processes was never proved rigorously. Presumably related, thermodynamic inequalities for semi-Markov processes remain elusive.

### v3

@@ IV Outlook
Recently, there has been a surge of interest in understanding and inferring irreversibility of coarse-grained observables. However, despite all efforts and advances, the picture is far from complete and many fundamental (and occasionally elementary) questions remain elusive.
Semi-Markov processes and milestoning. First, while the importance of semi-Markov processes (of order $n$) in stochastic thermodynamics is visibly growing, it seems to be underappreciated that parameterizing these requires much more effort and statistics. For example, whereas a Markov process is fully specified with all transition rates between states, a first-order semi-Markov process (i.e. renewal dynamics) requires for each given state the respective splitting probabilities and waiting-time distributions to transition into each of the neighboring states, respectively (and the waiting-time distributions require a parameterization themselves, say a sum of exponential functions). What is worse, a second-order semi-Markov process requires for each given state the splitting probabilities and waiting-time distributions to transition into each of the neighboring states conditioned on all incoming transitions, (and the waiting-time distributions require a parameterization themselves). Typically, we do not know a priori the order of the semi-Markov process unless we know both, the microscopic dynamics as well as the coarse graining. Moreover, the parameterization of semi-Markov processes becomes exponentially more challenging with the semi-Markov order. From a practical point of view one should therefore strive to approach renewal dynamics and hence invoke milestoning. However, formally milestoning has not yet been introduced to stochastic thermodynamics beyond renewal processes [62]. Moreover, we lack (sharp) thermodynamic inequalities, such as thermodynamic uncertainty relations and transport bounds, for milestoned dynamics. In particular, it would be interesting to also consider underdamped underlying dynamics. Another intriguing question, inspired by Ref. 67, concerns thermodynamic inference via milestoning of milestoned trajectories.
Thermodynamics of semi-Markov processes. Continuing with semi-Markov processes, rigorous results are only available for renewal dynamics. The results for the thermodynamic entropy production rate for general $n\geq 2$ semi-Markov processes was never proved rigorously. Presumably related, thermodynamic inequalities for semi-Markov processes remain elusive (with notable exceptions for the special case of first-order semi-Markov processes, see Ref. [123]).
@@
Appendix A The “Stratonovich term” is not a Stratonovich term
As mentioned in the main part, the last term in Eq. (8) is often mentioned to be a consequence of the “Stratonovich discretization scheme”, see, e.g., Ref. [54]. We stress that there is no such choice of convention, as the dynamics is already specified by the Langevin equation in Eq. (2) (which for additive noise does not even depend on the convention). In particular, since the dynamics are specified by Eq. (2), it should be clear that the action should be unique, which is why Eq. (8) is the only correct result for the action. The last term only (and obviously) disappears in the small-noise or low-temperature limit $D\to 0$, as considered in Freidlin-Wentzell theory [52].
To emphasize the uniqueness of the result for the action, we now list different ways to arrive at Eq. (8) which all (if performed correctly) give the same result. Moreover, recall that they also have to give the same result for the action in Eq. (8) since the statistics of the paths generated by Eq. (2) are the same for all conventions, because we only consider additive noise.
[…]
Another strategy to derive the action is to discretize the process in time, different discretization schemes virtually appear to correspond to different conventions. However, in the limit of continuous time, the different schemes all yield the action in Eq. (8), see Eqs. (17)-(30) in Ref. [133]. Note that Ref. [133] even shows the validity of Eq. (8) (and its failure without the last term) quantitatively for experimental data.
Another reference for Eq. (8) is Eq. (2.2b) in Ref. [134], where also in Eq. (2.2b) it is stated that the last term only vanishes in the small-noise or low-temperature limit, in agreement with Freidlin-Wentzell theory [52].